OpenStudy (anonymous):

integrate (1-square root(1+x))^2

OpenStudy (anonymous):

\[(1-\sqrt{1+x})^2 = 1 -2\sqrt{1+x} +1+x \]= \[\int\limits_{}^{} (2 +x ) dx - 2\int\limits_{}^{} \sqrt{1+x}\] for the second one use u = 1+x du = 1dx if u need more help please reply